OpenStudy (anonymous):

If my age is divided by 3, the remainder is 2. If my age is divided by 5, the remainder is also 2. If my age is divided by 7, the remainder is 5. How old am I?

OpenStudy (anonymous):

If my age is divided by 3, the remainder is 2. If my age is divided by 5, the remainder is also 2. If my age is divided by 7, the remainder is 5. How old am I?

OpenStudy (xapproachesinfinity):

we could set it this way a=3b+2 a=5c+2 a=7d+5 a= is your age b,c,d is what you multiply 3,5,7 to get close to your age with remainder 2,2,5 a-3b=2 a-5c=2 a-7d=5 solve this system you could have other different ways

OpenStudy (xapproachesinfinity):

actually this went get you any way! but it gives us a hint (a-2) is multiple of 5 and 3

OpenStudy (xapproachesinfinity):

common multiples of 5 and 3 {5,10,15,20,25,30.....} (3,6,9,12,15,18, 21,...} so least common multiple is 15 but we have another problem here a-5 is a multiple of 7 {7, 14, 21, 28, 35, 42, 49...}

OpenStudy (xapproachesinfinity):

so what we need now is to work multiple of 15 add 2 and multiple of 7 add 5 {17, 32, 47, 62...} for 15 {12, 19, 26, 33, 40. 47, 54...}' we observe that 47 is common here so it has to be the solution

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

That answer was in the back of my head

OpenStudy (xapproachesinfinity):

no checking! 47 divided by 3 we g3t remainder of 2 47 divided by 5 we get remainder of 2 47 divided by 7 we remainder of 5 so it is indeed 47

OpenStudy (xapproachesinfinity):

get*

OpenStudy (xapproachesinfinity):

you are welcome^_^

OpenStudy (anonymous):

I want to ask you another question Classify all the numbers that leave a remainder of 3 when divided by 5 and a remainder of 1 when divided by 3

OpenStudy (xapproachesinfinity):

hmmm x=5a+3 y=3b+1 so x is multiple of 5 add 3 and y is a multiple of 3 add 1 for 3 {5,10,15, 20,25,30...} {8,13,18, 23,28,33...} for 3 {3,6,12,15,18,21...} {4,7,13,16,19,22,25...}

OpenStudy (anonymous):

are u surre

OpenStudy (xapproachesinfinity):

that's not the answer we need to see what number have that condition btw all those numbers

OpenStudy (anonymous):

ohh

OpenStudy (xapproachesinfinity):

{8,13,18, 23,28,33...} {4,7,13,16,19,22,25...} you need to see it here

OpenStudy (xapproachesinfinity):

it seems that there is only one number 13 that satisfies this conditions however you need to extend those to braces you might find other numbers that have the same condition, i'm too lazy to continue counting this haha

OpenStudy (xapproachesinfinity):

13/5 remainder 3 13/3 remainder 1 you need to see it like this

OpenStudy (anonymous):

is there more cause they say its a list

OpenStudy (xapproachesinfinity):

could be! i only did few you will have to continue the list of two and choose the common ones and see if they satisfy the same condition

OpenStudy (anonymous):

28 works too

OpenStudy (xapproachesinfinity):

yeah! that one too

OpenStudy (xapproachesinfinity):

well 28 is common btw the two, if i continued i would get 28 after 25

OpenStudy (xapproachesinfinity):

they have to be common carry the on the rest tell you see some pattern or something you might get a conclusion

OpenStudy (anonymous):

can u try to find a equation

OpenStudy (xapproachesinfinity):

hmmm an equation, don't it works that way haha

OpenStudy (anonymous):

43 works and I found the equation and why

OpenStudy (xapproachesinfinity):

oh yeah! sounds good! let me see

OpenStudy (anonymous):

the numbers that works are 13 + 5 * 3

OpenStudy (anonymous):

then 28 + 5*3

OpenStudy (anonymous):

can u help me make an equation with variables

OpenStudy (xapproachesinfinity):

oh very clever my friend that is the pattern here

OpenStudy (anonymous):

thanks I am a human calculator

OpenStudy (anonymous):

what grade u in I am going to 8

OpenStudy (xapproachesinfinity):

i was too lazy to carry on hehe!

OpenStudy (anonymous):

please help me make an equation

OpenStudy (xapproachesinfinity):

oh this a sequence my friend the first term being 13 and the common difference of 15 have you studied this! about my grade, i'm a college student

OpenStudy (anonymous):

I did but this sequence can go on forever and I have to make a list

OpenStudy (xapproachesinfinity):

well have you studied sequences or not!

OpenStudy (anonymous):

i did

OpenStudy (xapproachesinfinity):

it doesn't matter if it goes for ever

OpenStudy (anonymous):

i am taking the shsat

OpenStudy (xapproachesinfinity):

ok now we can proceed! this sequence is what kind of sequences?

OpenStudy (xapproachesinfinity):

it is an arithmetic sequence, right? what do you know about this sequence

OpenStudy (xapproachesinfinity):

what is the general term

OpenStudy (anonymous):

it has a rate

OpenStudy (xapproachesinfinity):

not rate, we call it common different the general term is like this An=A1+(n-1)d A1 first term d common difference

OpenStudy (xapproachesinfinity):

do you remember this?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so the A1 is 13 and the comman difference is 15, right

OpenStudy (xapproachesinfinity):

yeah! so An=13+15(n-1) then we can say that the numbers that satisfy this equation are the numbers that satisfies the condition for any n>=1

OpenStudy (xapproachesinfinity):

actually we have the say that the term that that equation generates are the numbers that satisfy the condition for any n>=0 ====== my sentence above is incorrect!

OpenStudy (xapproachesinfinity):

actually we have to say that the terms that that equation generates are the numbers that satisfy the condition for any n>=0

OpenStudy (xapproachesinfinity):

sorry for bad spelling hehe

OpenStudy (anonymous):

so are we finished

OpenStudy (xapproachesinfinity):

yeah! that's all you need!

OpenStudy (xapproachesinfinity):

you don't have to make a list

OpenStudy (anonymous):

bye I will get your help next time

OpenStudy (xapproachesinfinity):

alright! perhaps i get of help to you! just tag me, and we'll see^_^

OpenStudy (anonymous):

ok

OpenStudy (xapproachesinfinity):

8 grade is high school? i don't understand the system here that much

OpenStudy (anonymous):

I am in middle school

OpenStudy (xapproachesinfinity):

i see, good luck^_^ study hard!